OpenStudy (anonymous):

The logarithm quotient is the sum of the logarithms of the numerator and the denominator.True or False?

OpenStudy (anonymous):

False

OpenStudy (anonymous):

\[\huge \log \frac{a}{b} = \log a- \log b\]

OpenStudy (anonymous):

False, dpasingh is right